Complete the following table and answer the accompanying questions Contral total Cost C( Benefits N(O) Benefit MC Marginal Ret BeneFit MNB MC 1,209 1,400 1,590 10e 101 102 103 104 105 106 107 108 109 950 210 60 80 1,940 2,100 2,250 2,390 2,520 2.640 2,750 100 110 120 130 140 150 166 110 a. At what level of the control variable are net benefits maximized? b. What is the relation between marginal benefit and marginal cost at this level of the variable? Marginal cost is slightly larger than marginal beneft. Marginal cost is much smaller than marginal benefit Marginal cost is much larger than marginal benefit Marginal cost is slightly smaller than marginal benefit.
Answer #1

Answer:

C(Q)=C(Q-1)+MC(Q)

N(Q)=B(Q)-C(Q)

MB(Q)=B(Q)-B(Q-1)

MNB(Q)=MB(Q)-MC(Q)

For Q=101

C=950+70=1020

N=1400-1020=380

MB=1400-1200=200

MNB=200-70=130

Q B C N MB MC MNB
100 1200 950 250 210 60 150
101 1400 1020 380 200 70 130
102 1590 1100 490 190 80 110
103 1770 1190 580 180 90 90
104 1940 1290 650 170 100 70
105 2100 1400 700 160 110 50
106 2250 1520 730 150 120 30
107 2390 1650 740 140 130 10
108 2520 1790 730 130 140 -10
109 2640 1940 700 120 150 -30
110 2750 2100 650 110 160 -50

:

A) From Above table we found net benefit maximize at Q=107 and Maximum Net benefit N=740

B) Option C:Marginal Cost is slightly smaller than marginal Benefit

Since Marginal cost at Q=107 is 130 and Marginal Benefit is 140